# Function-Correcting Codes for Insertion-Deletion Channels

### v1

@@ III-B Simplified Redundancy Lower Bounds
Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{2^{k}}\}$ be $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{M})$-code of length $N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}}))$ then $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{M}\}$ is a $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{M})$-code. Hence,
@@
From Theorem 1 we get $r^{f}_{ID}(k,t)\geq N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{M})).$
Since $E\geq 2$ and from [10132545][Corollary 1] we know that there exists $x,x^{\prime}\in\mathbb{F}_{2}^{k}$ with $d_{H}(x,x^{\prime})=1$ and $f(x)\neq f(x^{\prime})$ therefore using the inequality, $2\leq d_{ID}(x,x^{\prime})\leq 2d_{H}(x,x^{\prime})$ we get $d_{ID}(x,x^{\prime})=2$, that is there will always exist $x,x^{\prime}\in\mathbb{F}_{2}^{k}$ with $d_{ID}(x,x^{\prime})=2$ such that $f(x)\neq f(x^{\prime})$ whenever $E\geq 2.$ So, in particular for $M=2,$ we can say $r^{f}_{ID}(k,t)\geq N^{(1)}_{ID}(\bm{I}_{f}^{(2)}(2,2t-1)).$ Consider the following repetition code of length $t$, $\mathcal{C}=\{(0,0,\ldots,0),(1,1,\ldots,1)\}$ then $d_{ID}(\mathcal{C})=2t$ and $\mathcal{C}$ is a $N^{(1)}_{ID}(\bm{I}_{f}^{(2)}(2,2t-1))$ code hence $N^{(1)}_{ID}(\bm{I}_{f}^{(2)}(2,2t-1))=t.$
As pointed out in [10132545], finding optimal length of irregular distance codes over complete set of message vectors is quite difficult. In order to obtain a more computationally easier bound as compared to the Theorem 1 we define the concept of function distance and function distance matrices.
Definition 14 (Function Distance).
@@
$\scriptsize d_{ID}^{f}(f_{1},f_{2})=\min_{x_{1},x_{2}\in\mathbb{F}_{q}^{k}}d_{ID}(x_{1},x_{2})\quad\text{s.t.}\text{ }f(x_{1})=f_{1}\text{ and }f(x_{2})=f_{2}.$
Definition 15 (Function Distance Matrix).
The function insdel matrices for function $f$ are square matrices of order $E=|\text{Im}(f)|$ denoted by $\bm{I_{f}^{(1)}}(t,f_{1},\ldots,f_{E})$ and $\bm{I_{f}^{(2)}}(t,f_{1},\ldots,f_{E})$ respectively and whose entries are given by:
@@
Consider the following method of construction for FCIC. Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{E}\}$ be a set of codewords of length $r$ such that $r\geq k$ and $d_{ID}(p_{i},p_{j})\geq 2t+1+2k-d_{ID}^{f}(f_{i},f_{j}),\quad\forall i,j\in[E]$. Let $p_{i}$ be the redundancy vector corresponding to all $u\in\mathbb{F}_{2}^{k}$ such that $f(u)=f_{i}$, i.e., all message vectors mapping to the same function value have the same redundancy vector. From 8 it follows that for any $\mathbf{u}_{i},\mathbf{u}_{j}\in\mathbb{F}_{2}^{k}$ with $f(\mathbf{u}_{i})=f_{i}$, $f(\mathbf{u}_{j})=f_{j}$, $f_{i}\neq f_{j}$, we have
$d_{ID}(\mathrm{Enc}(\mathbf{u}_{i}),\mathrm{Enc}(\mathbf{u}_{j}))$ $\geq d_{ID}(\mathbf{u}_{i},\mathbf{u}_{j})+d_{ID}(p_{i},p_{j})$
$-2min(k,r)$
$\geq d_{ID}(u_{i},u_{j})+2t+1+2k$
$-d_{ID}^{f}(f_{i},f_{j})-2k$
$\geq 2t+1.$
According to the definition of FCIDC, we have constructed an FCIDC. Hence,
$r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}\big(I_{f}^{(2)}(t,f_{1},\ldots,f_{E});k\big).$
Deriving a general formula for the function distance $d_{ID}^{f}(f_{i},f_{j})$ is not always straightforward. Therefore, when an exact expression is unavailable, a lower bound on $d_{ID}^{f}(f_{i},f_{j})$ may be used to estimate an upper bound on the optimal redundancy.
Lemma 9.
Let $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a function with expressivness $E=Im(f)$ and let $\big\{a_{ij}\big\}_{1\leq i,j\leq E}$ be a set of $E$ non-negative integers such that $a_{ij}\leq d_{f}(f_{i},f_{j})$ for all $i,j$. Then,
$N_{ID}^{(2)}(I_{f}^{(2)}(t,f_{1},f_{2},...,f_{E});k)\leq N_{ID}^{(2)}(I;k)$
@@
Let $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{E}\}$ be a length $N_{ID}^{(2)}(I;k)$, irregular insdel distance code of type $2$ corresponding to matrix $I$. Then, for all $i\neq j$
@@
Hence, $\mathcal{C}$ is an irregular insdel distance code of type $2$ for function distance matrix $\bm{I_{f}^{(2)}}(t,f_{1},\ldots,f_{E})$. Therefore, $N_{ID}^{(2)}(I_{f}^{(2)}(t,f_{1},f_{2},...,f_{E});k)\leq N_{ID}^{(2)}(I;k).$
Combining Lemma 9 and Theorem 2 yields the following corollary.
Corollary 2.
Let $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a function with expressivness $E=Im(f)$ and let $\big\{a_{ij}\big\}_{1\leq i,j\leq E}$ be a set of $E$ non-negative integers such that $a_{ij}\leq d_{f}(f_{i},f_{j})$ for all $i,j$. Then,
@@
The next lemma is derived from the proof of [hayashi2020list, Lemma 1] by assuming the existence of a common binary super-sequence $v$ of length $N$ for all codewords in the binary insdel code, and setting $t_{D}=0$ (number of deletions) and $t_{I}=N-n$ (number of insertions), where $n$ is the length of the codewords.
Lemma 10.
[hayashi2020list] Let $C\subseteq\mathbb{F}_{2}^{r}$ be a code of minimum insdel distance $d$ and size $M$. Suppose that there exists $v\in\mathbb{F}_{2}^{N}$ that contains every codeword in $C$ as a subsequence, then
@@
Using the result from the previous lemma, we will now establish a lower bound on $N_{ID}^{(1)}(I)$ for a given distance matrix $I$. This bound can be viewed as a generalisation of the Plotkin-type bound for insdel codes with irregular distance requirements.
Lemma 11.
For $N\in\mathbb{N}$ and any matrix $I\in\mathbb{N}_{0}^{M\times M}$, such that $N\geq\frac{4S}{M^{2}}$,
@@
Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{n}\}$ be a length $r$, type $1$ irregular insdel distance code corresponding to matrix $I$ and let $v$ be an $N$-length binary sequence such that every codeword in $\mathcal{P}$ is a subsequence of $v$. Then from Lemma 10, $\sum_{i,j;i<j}d_{ID}(p_{i},p_{j})\leq\frac{M^{2}(N-r)r}{N}$.Since,
@@
Corollary 3.
For $N\in\mathbb{N}$ and any matrix $I\in\mathbb{N}_{0}^{M\times M}$, such that $N=\frac{4\sum_{i,j:i<j}[I]_{ij}}{M^{2}}$,
@@
The result in Corollary 3 can also be derived from the plotkin-like bound for codes with irregular distance requirement given in [10132545] for the Hamming metric, together with the observation that if the insdel distance between two codewords is $d$, then their Hamming distance is at least $\frac{d}{2}$. For completeness, we present below a lemma that formally proves this result using the aforementioned approach.
Lemma 12.
@@
For any $i<j$, we have from Lemma 1
$d_{H}(p_{i},p_{j})\;\geq\;\tfrac{1}{2}\,d_{\mathrm{ID}}(p_{i},p_{j})\;\geq\;\tfrac{1}{2}\,[I]_{ij},$
@@
From Lemma 5 we know in case of even $M$
@@
$N^{(1)}_{\mathrm{ID}}(\bm{I})\leq\min_{r\in\mathbb{N}}\left\{r\ \middle|\ 2^{r}>\max_{j\in\{1,2,\ldots,M\}}\sum_{i=1}^{j-1}\left|B_{E}\big(r,[\bm{I}]_{\pi(i)\pi(j)}-1\big)\right|\right\},$
@@
$N^{(2)}_{\mathrm{ID}}(\bm{I};K)\leq\min_{r\geq K}\left\{r\ \middle|\ 2^{r}>\max_{j\in\{1,2,\ldots,M\}}\sum_{i=1}^{j-1}\left|B_{E}\big(r,[\bm{I}]_{\pi(i)\pi(j)}-1\big)\right|\right\},$
@@
$B_{ID}(r,t)=\{\,y\in\mathbb{F}_{2}^{r}\ \big|\ d_{\mathrm{ID}}(x,y)\leq t\,\}$
denotes the editing ball of radius $t$ centered at a vector $x$ in the insdel metric.
Proof.
We present an iterative way of selecting valid codewords for constructing an irregular insdel distance code of type 2 (type 1) for any distance matrix $\bm{I}$. For simplicity, first choose $\pi$ as the identity permutation. Select any arbitrary vector ${p_{1}}\in\mathbb{F}_{2}^{r}$ as the first codeword of the irregular insdel distance code of type 2 (type 1) such that $r\geq K(r\in\mathbb{N})$. For the second codeword $p_{2}$}, the distance requirement
$d_{\text{ID}}({p_{1}},{p_{2}})\geq[\bm{I}]_{12}$
must be satisfied. The existence of such a $p_{2}$} is guaranteed if
$2^{r}>\big|B_{ID}\big({p_{1}},[\bm{I}]_{12}-1\big)\big|,$
Next, choose the third codeword $p_{3}$} such that
$d_{\text{ID}}({p_{1}},{p_{3}})\geq[\bm{I}]_{13}\quad\text{and}\quad d_{\text{ID}}({p_{2}},{p_{3}})\geq[\bm{I}]_{23}.$
For $p_{3}$} to exist, it suffices that
$2^{r}>\big|B_{ID}\big({p_{1}},[\bm{I}]_{13}-1\big)\big|+\big|B_{ID}\big({p_{2}},[\bm{I}]_{23}-1\big)\big|.$
Proceeding inductively, at the $j$-th step we must choose $p_{j}$} so that for all $i<j$,
$d_{\text{ID}}({p_{i}},{p_{j}})\geq[\bm{I}]_{ij}.$
A sufficient condition for the existence of $p_{j}$} is
$2^{r}>\sum_{i=1}^{j-1}\big|B_{ID}\big({p_{i}},[\bm{I}]_{ij}-1\big)\big|.$
Since the choice of codewords can be made in any order, the same argument applies for any permutation $\pi$, which proves the lemma. ∎
IV VT-syndrome function

### v2

@@ 3.2 Simplified Redundancy Lower Bounds
Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{2^{k}}\}$ be $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}})$-code of length $N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}}))$ then $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{M}\}$ is a $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{M})$-code. Hence,
@@
From Theorem 3.17 we get $r^{f}_{ID}(k,t)\geq N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{M})).$
Since $E\geq 2$ for a non-constant function and from [1, Corollary 1] we know that there exists $x,x^{\prime}\in\mathbb{F}_{2}^{k}$ with $d_{H}(x,x^{\prime})=1$ and $f(x)\neq f(x^{\prime})$ therefore using the inequality, $2\leq d_{ID}(x,x^{\prime})\leq 2d_{H}(x,x^{\prime})$ we get $d_{ID}(x,x^{\prime})=2$, that is there will always exist $x,x^{\prime}\in\mathbb{F}_{2}^{k}$ with $d_{ID}(x,x^{\prime})=2$ such that $f(x)\neq f(x^{\prime})$ whenever $E\geq 2.$ So, in particular for $M=2,$ we can say $r^{f}_{ID}(k,t)\geq N^{(1)}_{ID}(2,2t).$ Consider the following repetition code of length $t$, $\mathcal{C}=\{(0,0,\ldots,0),(1,1,\ldots,1)\}$ then $d_{ID}(\mathcal{C})=2t$. Hence, $N^{(1)}_{ID}(2,2t)=t.$
As pointed out in [1], finding the optimal length of irregular distance codes over a complete set of message vectors is quite difficult. In order to obtain a more computationally easier bound as compared to the Theorem 3.17 we define the concept of function distance and function distance matrices.
Definition 3.20 (Function Distance).
@@
$d_{ID}^{f}(f_{1},f_{2})=\min_{x_{1},x_{2}\in\mathbb{F}_{2}^{k}}d_{ID}(x_{1},x_{2})\quad\text{s.t.}\text{ }f(x_{1})=f_{1}\text{ and }f(x_{2})=f_{2}.$
Definition 3.21 (Function Distance Matrices).
The function insdel-distance matrices for function $f$ are square matrices of order $E=|\text{Im}(f)|$ denoted by $\bm{I}_{f}^{(1)}(t,f_{1},\ldots,f_{E})$ and $\bm{I}_{f}^{(2)}(t,f_{1},\ldots,f_{E})$ respectively and whose entries are given by:
@@
Consider the following construction method for FCIDCs. Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{E}\}$ be a set of codewords of length $r$ such that $r\geq k$ and $d_{ID}(p_{i},p_{j})\geq 2t+2+2k-d_{ID}^{f}(f_{i},f_{j}),\quad\forall i,j\in[E]$. Let $p_{i}$ be the redundancy vector corresponding to all $u\in\mathbb{F}_{2}^{k}$ such that $f(u)=f_{i}$, i.e., all message vectors mapping to the same function value have the same redundancy vector. From 3.14 it follows that for any $\mathbf{u}_{i},\mathbf{u}_{j}\in\mathbb{F}_{2}^{k}$ with $f(\mathbf{u}_{i})=f_{i}$, $f(\mathbf{u}_{j})=f_{j}$, $f_{i}\neq f_{j}$, we have
$d_{ID}(\psi(\mathbf{u}_{i}),\psi(\mathbf{u}_{j}))$ $\geq d_{ID}(\mathbf{u}_{i},\mathbf{u}_{j})+d_{ID}(p_{i},p_{j})-2\cdot min\{k,r\}$
$\geq d_{ID}(u_{i},u_{j})+2t+2+2k-d_{ID}^{f}(f_{i},f_{j})-2k$
$\geq 2t+2.$
According to the definition of FCIDCs, we have constructed an FCIDC. Hence,
$r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}\big({\bm{I}}_{f}^{(2)}(t,f_{1},\ldots,f_{E});k\big).$
It is not always easy to derive a generic formula for the function distance $d_{ID}^{f}(f_{i},f_{j})$. Consequently, an upper bound on the optimal redundancy can be estimated using a lower constraint on $d_{ID}^{f}(f_{i},f_{j})$ in the absence of an explicit equation.
Lemma 3.23.
Let $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a function with expressivness $E=|Im(f)|$ and let $\big\{a_{ij}\big\}_{1\leq i,j\leq E}$ be a set of $E$ non-negative integers such that $a_{ij}\leq d_{ID}^{f}(f_{i},f_{j})$ for all $i,j$. Then,
$N_{ID}^{(2)}(\bm{I}_{f}^{(2)}(t,f_{1},f_{2},...,f_{E});k)\leq N_{ID}^{(2)}(I;k)$
@@
Let $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{E}\}$ be a length $N_{ID}^{(2)}(I;k)$, irregular insdel-distance code of type $2$ corresponding to matrix $I$. Then, for all $i\neq j$
@@
Hence, $\mathcal{C}$ is an irregular insdel-distance code of type $2$ for function distance matrix ${\bm{I}_{f}^{(2)}}(t,f_{1},\ldots,f_{E})$. Therefore, $N_{ID}^{(2)}(\bm{I}_{f}^{(2)}(t,f_{1},f_{2},...,f_{E});k)\leq N_{ID}^{(2)}(I;k).$
Combining Lemma 3.23 and Theorem 3.22 yields the following corollary.
Corollary 3.24.
Let $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a function with expressivness $E=|Im(f)|$ and let $\big\{a_{ij}\big\}_{1\leq i,j\leq E}$ be a set of $E$ non-negative integers such that $a_{ij}\leq d_{f}(f_{i},f_{j})$ for all $i,j$. Then,
@@
The next lemma is derived from the proof of [17, Lemma 1] by assuming the existence of a common binary super-sequence $v$ of length $N$ for all codewords in the binary insdel code, and setting $t_{D}=0$ (number of deletions) and $t_{I}=N-n$ (number of insertions), where $n$ is the length of the codewords.
Lemma 3.25.
[17] Let $C\subseteq\mathbb{F}_{2}^{r}$ be a code of minimum insdel distance $d$ and size $M$. Suppose that there exists $v\in\mathbb{F}_{2}^{N}$ that contains every codeword in $C$ as a subsequence, then
@@
Using the result from the previous lemma, we will now establish a lower bound on $N_{ID}^{(1)}(I)$ for a given distance matrix $I$. This bound can be viewed as a generalization of the Plotkin-type bound for insdel codes with irregular distance requirements.
Lemma 3.26.
For $N\in\mathbb{N}$ and any matrix $\bm{I}\in\mathbb{N}_{0}^{M\times M}$, such that $N\geq\frac{4S}{M^{2}}$,
@@
Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{M}\}$ be a length $r$, type $1$ irregular insdel-distance code corresponding to matrix $I$ and let $v$ be an $N$-length binary sequence such that every codeword in $\mathcal{P}$ is a subsequence of $v$. Then from Lemma 3.25, $\sum_{i,j;i<j}d_{ID}(p_{i},p_{j})\leq\frac{M^{2}(N-r)r}{N}$. Since,
@@
Corollary 3.27.
For $N\in\mathbb{N}$ and any matrix $\bm{I}\in\mathbb{N}_{0}^{M\times M}$, such that $N=\frac{4\sum_{i,j:i<j}[I]_{ij}}{M^{2}}$,
@@
The result in Corollary 3.27 can also be derived from the plotkin-like bound for codes with irregular distance requirement given in [1] for the Hamming metric, together with the observation that if the insdel distance between two codewords is $d$, then their Hamming distance is at least $\frac{d}{2}$. For completeness, we present below a lemma that formally proves this result using the aforementioned approach.
Lemma 3.28.
@@
For any $i<j$, we have from Lemma 2.8
$d_{H}(p_{i},p_{j})\;\geq\;\tfrac{1}{2}\,d_{ID}(p_{i},p_{j})\;\geq\;\tfrac{1}{2}\,[I]_{ij},$
@@
From Lemma 2.16, we know in case of even $M$
@@
$N^{(1)}_{ID}(\bm{I})\leq\min_{r\in\mathbb{N}}\left\{r\ \middle|\ 2^{r}>\max_{j\in\{1,2,\ldots,M\}}\sum_{i=1}^{j-1}B^{\max}_{ID}\big(r,[\bm{I}]_{\pi(i)\pi(j)}-2\big)\right\},$
@@
$N^{(2)}_{ID}(\bm{I};K)\leq\min_{r\geq K}\left\{r\ \middle|\ 2^{r}>\max_{j\in\{1,2,\ldots,M\}}\sum_{i=1}^{j-1}B^{\max}_{ID}\big(r,[\bm{I}]_{\pi(i)\pi(j)}-2\big)\right\}.$
$B^{\max}_{ID}(r,t)\triangleq\max_{x\in\mathbb{F}_{2}^{r}}\left|B_{ID}(x,t)\right|,\qquad B_{ID}(x,t)=\{\,y\in\mathbb{F}_{2}^{r}\mid d_{ID}(x,y)\leq t\,\}.$
Proof.
We present an iterative way of selecting valid codewords for constructing an irregular insdel-distance code of type 2 (type 1) for any distance matrix $\bm{I}$. For simplicity, first choose $\pi$ as the identity permutation. Select an arbitrary vector $p_{1}\in\mathbb{F}_{2}^{r}$ as the first codeword, where $r\geq K$ (or $r\in\mathbb{N}$ for type 1). For the second codeword $p_{2}$, the distance requirement
$d_{ID}(p_{1},p_{2})\geq[\bm{I}]_{12}$
must be satisfied. Such a $p_{2}$ exists provided
$2^{r}>\left|B_{ID}\big(p_{1},[\bm{I}]_{12}-2\big)\right|.$
Since the size of an insdel ball depends on its center, we upper bound it by the worst-case ball size $B^{\max}_{ID}(r,[\bm{I}]_{12}-2)$.
Next, choose the third codeword $p_{3}$ such that
$d_{ID}(p_{1},p_{3})\geq[\bm{I}]_{13}\quad\text{and}\quad d_{ID}(p_{2},p_{3})\geq[\bm{I}]_{23}.$
A sufficient condition for the existence of such a $p_{3}$ is
$2^{r}>\left|B_{ID}\big(p_{1},[\bm{I}]_{13}-2\big)\right|+\left|B_{ID}\big(p_{2},[\bm{I}]_{23}-2\big)\right|$
which is guaranteed if
$2^{r}>B^{\max}_{ID}\big(r,[\bm{I}]_{13}-2\big)+B^{\max}_{ID}\big(r,[\bm{I}]_{23}-2\big).$
Proceeding inductively, at the $j$-th step we must choose $p_{j}$ such that
$d_{ID}(p_{i},p_{j})\geq[\bm{I}]_{ij}\quad\text{for all }i<j.$
A sufficient condition for the existence of $p_{j}$ is
$2^{r}>\sum_{i=1}^{j-1}B^{\max}_{ID}\big(r,[\bm{I}]_{ij}-2\big).$
Since the codewords can be selected in any order, the same argument holds for any permutation $\pi$, completing the proof. ∎
Theorem 3.30.
For a positive integer $M$, an even positive integer $d$, and $K\geq 2$, we have
$N_{ID}^{(2)}(M,d;K)\leq\Bigg\lceil\frac{\ln M+(d-2)\ln\!\Big(\frac{2eK}{d-2}\Big)}{\ln 2}\Bigg\rceil.$
Proof.
By Lemma 3.29, the minimum length of a type 2 irregular insdel-distance code with $M$ codewords and distance requirement $d$ satisfies
$N^{(2)}_{ID}(M,d;K)$ $\leq\min_{r\geq K}\left\{r\,\middle|\,2^{r}>\max_{j\in\{1,\ldots,M\}}\sum_{i=1}^{j-1}B^{\max}_{ID}(r,d-2)\right\}$
$\leq\min_{r\geq K}\left\{r\,\middle|\,2^{r}>M\,B^{\max}_{ID}(r,d-2)\right\},$
@@
$B^{\max}_{ID}(r,t)\triangleq\max_{x\in\mathbb{F}_{2}^{r}}\left|B_{ID}(x,t)\right|.$
Using the upper bound on the insdel ball size from Lemma 2.11, we have
$B^{\max}_{ID}(r,d-2)\leq\frac{e^{d-2}\,r^{\frac{d-2}{2}}\,\left(r+\frac{d-2}{2}\right)^{\frac{d-2}{2}}}{\left(\frac{d-2}{2}\right)^{d-2}}.$
Let $d^{\prime}=d-2$. A sufficient condition for $M\cdot\,B^{\max}_{ID}(r,d^{\prime})<2^{r}$ is
$\frac{e^{d^{\prime}}\,r^{\frac{d^{\prime}}{2}}\,\left(r+\frac{d^{\prime}}{2}\right)^{\frac{d^{\prime}}{2}}}{\left(\frac{d^{\prime}}{2}\right)^{d^{\prime}}}\,M<2^{r}.$
Taking natural logarithms on both sides yields
$d^{\prime}+\frac{d^{\prime}}{2}\ln r+\frac{d^{\prime}}{2}\ln\!\left(r+\frac{d^{\prime}}{2}\right)-d^{\prime}\ln\!\left(\frac{d^{\prime}}{2}\right)+\ln M<r\ln 2.$
Using $d^{\prime}\ln\!\left(\frac{d^{\prime}}{2}\right)=-d^{\prime}\ln d^{\prime}+d^{\prime}\ln 2$ and the inequality $\ln\!\left(r+\frac{d^{\prime}}{2}\right)\geq\ln r,$ we obtain
$\ln M+d^{\prime}-d^{\prime}\ln d^{\prime}+d^{\prime}\ln 2<r\ln 2-d^{\prime}\ln r.$
Since $r\geq K$, we have $\ln r\geq\ln K$, and hence
$r\ln 2>\ln M+d^{\prime}+d^{\prime}\ln\!\Big(\frac{2K}{d^{\prime}}\Big).$
Dividing both sides by $\ln 2$ gives
$r>\frac{\ln M+d^{\prime}\ln\!\Big(\frac{2eK}{d^{\prime}}\Big)}{\ln 2}.$
Therefore,
$N_{ID}^{(2)}(M,d;K)\leq\Bigg\lceil\frac{\ln M+(d-2)\ln\!\Big(\frac{2eK}{d-2}\Big)}{\ln 2}\Bigg\rceil,$
which completes the proof. ∎
To bound $N_{ID}^{(2)}(\mathbf{I};K)$, we first relate it to $N_{ID}^{(2)}(M,I;K)$. The bound on the latter, provided by Theorem 3.30, then directly implies a bound on the former.
Corollary 3.31.
Let $\mathbf{I}\in\mathbb{N}_{0}^{M\times M}$, and let $I_{max}$ denote its largest entry. If $I_{max}\leq I$, then
$N_{ID}^{(2)}(\mathbf{I};K)\leq N_{ID}^{(2)}(M,I;K).$
4 VT-syndrome function